Sodium and copper have work function 2.3 eve and 4.5 ev. Then the ratio of the wavelength is nearest to
 

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

I suppose the question is asking about the ratio of cut-off wavelengths of the two metals.

We know thatW_0=h\nu=\dfrac{hc}{\lambda}
HenceW_0\propto \dfrac{1}{\lambda}
\implies \dfrac{\lambda_1}{\lambda_2}=\dfrac{4.5}{2.3}=1.96